OpenStudy (anonymous):

PLZ PLZHELP ILL GIVE MEDAL AND ILL WRITE A RECOMMENDATION LETTER AT THE WEBSITE OPENSTUDY LIFESAVERS PLZ PLZ HELP

OpenStudy (anonymous):

I can try to help, what is it

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

What is the first step in solving the equation three-tenths equals x divided by three plus one-tenth for x? Add one-tenth to both sides of the equal sign. Add three-tenths to both sides of the equal sign. Subtract one-tenth from both sides of the equal sign. Subtract three-tenths from both sides of the equal sign.

OpenStudy (anonymous):

\[\large \frac{3}{10}=\frac{x}{3}+\frac{1}{10}\] like that?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Use the process of elimination, what makes the most sense?

OpenStudy (anonymous):

if so, first step is to subtract \(\frac{1}{10}\) from both sides to get \[\large\frac{2}{10}=\frac{x}{3}\]

OpenStudy (andrewkaiser333):

it is a

OpenStudy (anonymous):

wait wait

OpenStudy (anonymous):

actually it is C, because it is always C

OpenStudy (andrewkaiser333):

subtract 1/10 from the one side to the other

OpenStudy (andrewkaiser333):

that is how it is done the algebric way

OpenStudy (andrewkaiser333):

i mean c

OpenStudy (anonymous):

How could you eliminate the fractions before proceeding to solve this equation? five-eighths equals three-eighths times t plus one-eighth Add one-eighth to both sides of the equation. Subtract one-eighth from both sides of the equation. Divide both sides of the equation by 8. Multiply both sides of the equation by 8.

OpenStudy (andrewkaiser333):

sorry

OpenStudy (andrewkaiser333):

i messed up

OpenStudy (andrewkaiser333):

i am tired i am going to get off bye

OpenStudy (andrewkaiser333):

sattelite is right and i would half id i typed it right sorry bout that

OpenStudy (anonymous):

How could you eliminate the fractions before proceeding to solve this equation? five-eighths equals three-eighths times t plus one-eighth Add one-eighth to both sides of the equation. Subtract one-eighth from both sides of the equation. Divide both sides of the equation by 8. Multiply both sides of the equation by 8.

OpenStudy (anonymous):

How could you eliminate the fractions before proceeding to solve this equation? five-eighths equals three-eighths times t plus one-eighth Add one-eighth to both sides of the equation. Subtract one-eighth from both sides of the equation. Divide both sides of the equation by 8. Multiply both sides of the equation by 8.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

plz someone help

OpenStudy (anonymous):